OpenStudy (anonymous):

What is the solution of the equation? 2 = 6p - 8 - 5p A. –10 B. –6 C. 2 D. 10

OpenStudy (anonymous):

.add and combine like terms

OpenStudy (anonymous):

2=6p-8-5p 2=p-8 -2 -2. P= -10